pooja195 (pooja195):

Brenda's school is selling tickets to a spring musical. On the first day of ticket sales the school sold 3 senior citizen tickets and 9 child tickets for a total of $75. The school took in $67 on the second day by selling 8 senior citizen tickets and 5 child tickets. What is the price each of one senior citizen ticket and one child ticket? @iambatman @iGreen @e.mccormick

OpenStudy (e.mccormick):

OK, so, how do you think you need to set it up? Obviously you need to make two equations to find the two unknowns.

pooja195 (pooja195):

Right thats what i dont know how to do I can't solve these without a verbal model

OpenStudy (e.mccormick):

So, lets take it one line at a time. On the first day of ticket sales the school sold 3 senior citizen tickets and 9 child tickets for a total of $75. So what numbers and what variables do you see there?

pooja195 (pooja195):

3 senior tickets 9 child tickets total $75

OpenStudy (e.mccormick):

Good. So, if we just shorten the names we can say: 3s 9c total $75 Now, can you put that into an equation?

pooja195 (pooja195):

3s+9c=75?

OpenStudy (e.mccormick):

Yep! Exactly! Now, do the same basic thing for " The school took in $67 on the second day by selling 8 senior citizen tickets and 5 child tickets."

pooja195 (pooja195):

8c+5c=67?

OpenStudy (e.mccormick):

oops

pooja195 (pooja195):

wait no 8s+5c=67?

OpenStudy (e.mccormick):

Not both c.

OpenStudy (e.mccormick):

OK, now you have: 3s+9c=75 8s+5c=67

pooja195 (pooja195):

so now all i do is solve right?

OpenStudy (e.mccormick):

Yep!

pooja195 (pooja195):

ok i have it from here thanks! :D

pooja195 (pooja195):

senior citizen ticket: $4, child ticket: $7 is this right?

OpenStudy (e.mccormick):

Yah.

OpenStudy (e.mccormick):

I'll show some other ways to solve it for the fun of it...

pooja195 (pooja195):

ok :)

OpenStudy (e.mccormick):

3s+9c=75 8s+5c=67 Lets see... I see three easy solutions to that. First, graphing: https://www.desmos.com/calculator/laerywuegn Now for some elimination in a matrix! \(\left[\begin{array}{cc|c} 3 & 9 & 75\\ 8 & 5 & 67\\ \end{array}\right]\) divide row one by 3. \(\left[\begin{array}{cc|c} 1 & 3 & 25\\ 8 & 5 & 67\\ \end{array}\right]\) Multiply row 1 by -8 and add to row 2. \(\left[\begin{array}{cc|c} 1 & 3 & 25\\ 0 & -19 & -133\\ \end{array}\right]\) Divide row 2 by -19. \(\left[\begin{array}{cc|c} 1 & 3 & 25\\ 0 & 1 & 7\\ \end{array}\right]\) Multiply row 2 by -3 and add to row 1: \(\left[\begin{array}{cc|c} 1 & 0 & 4\\ 0 & 1 & 7\\ \end{array}\right]\) Presto! And then there is basic substitution 3s+9c=75 8s+5c=67 3s=75-9c s=25-3c So: 8(25-3c)+5c=67 200-24c+5c=67 -19c=67-200 -19c=-133 c=7 3s+9(7)=75 3s+63=75 3s=12 s=4

OpenStudy (e.mccormick):

Took me more time to type up the augmented matrix than to do the math. LOL

pooja195 (pooja195):

O_O wow theres so many ways to do one problem O_O i used elimination but i think the substituition way its alot more easier :) you make it look like a picece of cake xD

OpenStudy (e.mccormick):

Well, when I saw that all the parts of the first equation were dividible by 3...

pooja195 (pooja195):

Ah ok that makes sense that would've been easier to do XD thanks :D